OpenStudy (anonymous):

Help please! Really easy question! Giving medals! What is 7 1/4 x 2?

OpenStudy (anonymous):

Do you mean 7=1/4x*2 If so x=14

OpenStudy (anonymous):

\[7\frac{ 1 }{ 4 } \times 2\]?

OpenStudy (anonymous):

no I meant what does 7 and 1/4 x 2 equal

OpenStudy (anonymous):

15

OpenStudy (anonymous):

@Johnbc right

OpenStudy (anonymous):

The answer is then 15 like I said

OpenStudy (anonymous):

ok then! thanks